OpenStudy (anonymous):

For the quadratic function, find the x-intercepts and the y-intercept: m(x) = 10x^2 - 17x + 3

OpenStudy (raden):

for x-intercepts, take m(x) = 0 or 10x^2 - 17x + 3 = 0, solve for x for y-intercept, take x = 0

OpenStudy (anonymous):

So I have to factor it out? And then solve for 0?

OpenStudy (raden):

yes... go ahead

OpenStudy (anonymous):

But there is no way to factor this out so I use the quadratic formula?

OpenStudy (raden):

10x^2 - 17x + 3 = (2x-3)(5x-1)

OpenStudy (raden):

now find zeros all factors, take 2x-3 = 0 and 5x-1 = 0, solve for both

OpenStudy (anonymous):

Wait how did you factor the equation out?

OpenStudy (raden):

ok, for 10x^2 - 17x + 3 first, we find factor of 10x^2, it can be 2x times 5x or reverse (sometimes satisfy with x times 10x, but look situation for middle term) next, find factor of 3, it can be 1 times 3 or reverse) for ilustration, look here :